OpenStudy (anonymous):

Find the sun of the series 6+12+24+48+....+1536 show work so I can understand and don't ask questions because it won't let me reply to you all.

OpenStudy (zzr0ck3r):

What are you having a problem with?

OpenStudy (anitaasks):

This is what I got \[\sum_{k=1}^{256} 6k\] = \[6\sum_{k=1}^{256}k\] => use formula => 6(n(n+1)/2) = 6(256(257)/2) = 197376

OpenStudy (zzr0ck3r):

I think its more like \(\large\sum_{n=1}^86*2^n=\large6\sum_{n=1}^82^n\) @anitaasks maybe you can help him from here